OpenStudy (anonymous):

in a circle of coordinates for the endpoints of a diameter are (-4,2) amd (2,-4). find the coordinates of the center of the circle. Help please!! thank you

OpenStudy (anonymous):

Just find the average value between the points. Do you know how to do that?

OpenStudy (anonymous):

no i dont :(

OpenStudy (anonymous):

Given two points \((x_1,y_1)\) and \((x_2,y_2)\), the average point is given as follows.\[\left(\frac{x_2+x_1}{2},\frac{y_2+y_1}{2}\right)\] This is the same as finding the average \(x\)-coordiante and the average \(y\)-coordinate if you prefer to think of it that way.

OpenStudy (anonymous):

ooo

OpenStudy (anonymous):

so it would be x2+x1? I thought it would be like x1+y1 and the other one would be x2+y2

OpenStudy (anonymous):

It is \(x_2+x_1\). Treat each of the axes separately.

OpenStudy (anonymous):

oook